5 women and 6 men are on a softball team. Two of them are brothers. How many ways can you form a committee of 2 women and 3 men if you keep the brothers together, either both on the committee or both off the committee? What is the probability that the committee will have 2 women and 3 men and that the brothers will be together, either both on the committee or both off the committee?
I tried 5C3 x 5C2 for the first part and (5C3 x 5C2)11C5 for the second part but I don't think it's right. Any help would be greatly appreciated.
Thanks

Answer by sudhanshu_kmr(1152)


There are 5 women and 6 men. We have to select 2 women and 3 men.
First case: both brothers are in committee
as both are in committee, now we have to choose 1 from remaining 4 men
no. of ways = 5C2 * 4C1 = 10 * 4 = 40
Second case : both are not in committee
now we have to choose all 3 men from remaining 4
no. of ways = 5C2 * 4C3 = 10 * 4 = 40
total no. of ways = 40 + 40 = 80
Your question related to probability is not cleared...
please resubmit